ABCD is an isosceles trapezoid (named counterclockwise) with median XY and diagonals AC and BD. If the measurement of angle DAB=2x-10 and the measurement of angle ADC is 3x+40, what is the measurement lf angle CBA
Answer by ikleyn(53937)

.
First,  angle A plus angle D is 180 degrees,  or


    (2x-10) + (3x + 40) = 180

           5x           = 180 + 10 - 40 = 150

            x           = 150/5 = 30,

Hence,  angle A is 2x-10 = 2*30-10 = 50  degrees.


Angle B (which is CBA) is congruent to angle A;  so, angle CBA is 50 degrees.    ANSWER

Solved.

The median XY is irrelevant to the problem.

The diagonals AC and BD are irrelevant, too.